# Self-Introspective Decoding: Alleviating Hallucinations for Large Vision-Language Models

### v2

@@ A.6 More ABLATION ANALYSES
Quantitative Attention Scores Analyses. Here we quantitatively analyze vision attention score in terms of POPE Li et al. (2023d) and CHAIR Rohrbach et al. (2018) metrics. Concretely, we select the top-100 and least-100 important vision tokens out of a total of 576 vision tokens of LLaVA-1.5 7B based on attention score based on Eq. 5. Visual and Instruction Disturbance (VD and ID) are also employed as inputs for analyses. Quantitative results in Table 8 illustrate that 100 out of 576 vision tokens with high attention scores greatly maintain original ability, while low attention score tokens reach almost 50 accuracy for the binary classification problem, which indicates attention scores are a good indicator for vision token importance. As for VD and ID, disturbance in raw input does not obviously harm the LVLMs’ discrimination ability, as indicated by the POPE metric. However, VD and ID significantly compromise the open-end generation tasks reflected by the CHAIR metric (LVLMs tend to refuse to ID as shown in Fig. 5 and 10).
@@
Token Selection Strategies Analyses. To validate the effectiveness of SID in selecting low attention scores to induce vision-and-text association hallucination, we further conduct quantitative experiments under different vision token selection strategies with the same preserved vision token number and Layer =3 as ours. Table 9 shows that vision tokens with high attention scores degrade obviously, as it does not amplify contextual hallucinations rather than retain original important information. Contrastive decoding does not benefit from subtracting hallucinations amplified by the disturbed inputs rather than suffers from the potential disturbance noise. Selecting random vision tokens brings improvements in the adversarial setting because randomly selected vision tokens amplify the over-reliance on statistical bias and language priors, similar to Vision CD (Leng et al., 2024) and Instruction CD (Wang et al., 2024a). However, token-level random disturbance also induces uncertainty noise, resulting in the inferior performance in the random setting to greedy decoding. Overall, these experiments further validate the rationality of our token selection strategy based on attention sores.
@@
Larger-scale LVLM Backbone.

### v3

@@ A.7 More Ablation Analyses
Quantitative Attention Scores Analyses. Here we quantitatively analyze vision attention score in terms of POPE (Li et al., 2023d) and CHAIR (Rohrbach et al., 2018) metrics. Concretely, we select the top-100 and least-100 important vision tokens out of a total of 576 vision tokens of LLaVA-1.5 7B based on attention score based on Eq. 5. Visual and Instruction Disturbance (VD and ID) are also employed as inputs for analyses. Quantitative results in Table 12 illustrate that 100 out of 576 vision tokens with high attention scores greatly maintain original ability, while low attention score tokens reach almost 50 accuracy for the binary classification problem, which indicates attention scores are a good indicator for vision token importance. As for VD and ID, disturbance in raw input does not obviously harm the LVLMs’ discrimination ability, as indicated by the POPE metric. However, VD and ID significantly compromise the open-end generation tasks reflected by the CHAIR metric (LVLMs tend to refuse to ID as shown in Fig. 5 and 10).
@@
Token Selection Strategies Analyses. To validate the effectiveness of SID in selecting low attention scores to induce vision-and-text association hallucination, we further conduct quantitative experiments under different vision token selection strategies with the same preserved vision token number and Layer =3 as ours. Table 13 shows that vision tokens with high attention scores degrade obviously, as it does not amplify contextual hallucinations rather than retain original important information. Contrastive decoding does not benefit from subtracting hallucinations amplified by the disturbed inputs rather than suffers from the potential disturbance noise. Selecting random vision tokens brings improvements in the adversarial setting because randomly selected vision tokens amplify the over-reliance on statistical bias and language priors, similar to Vision CD (Leng et al., 2024) and Instruction CD (Wang et al., 2024a). However, token-level random disturbance also induces uncertainty noise, resulting in the inferior performance in the random setting to greedy decoding. Moreover, AVISC (Woo et al., 2024), in contrast to ours, preserves outlier high attention tokens (named ’blind token’) and substracts output logits to counteract the overemphasis of ’blind token.’ In this way, AVISC promotes balanced consideration of all tokens to alleviate hallucinations. However, Table 13 illustrates that Top-100 vision tokens with high attention scores can largely maintain the original performance. ’blind token’ tends to have a high probability of target class logits, and contrastive decoding does not improve the target class’s probability while might bring extra noise. Table 13 indicates AVISC still degrades the greedy decoding to some extent, which indicates the attentional vision re-calibration strategy of AVISC induces some annoying noise. Overall, these experiments further validate the rationality of our token selection strategy based on attention sores.
@@
| AVISC | 88.4 | 88.1 | 79.8 | 80.5 | 45.3 | 14.7 |
| Sampling | 84.9 | 83.2 | 78.7 | 78.9 | 51.3 | 16.8 |
[…]
| AVISC | 87.9 | 87.9 | 77.5 | 79.6 | 46.6 | 12.5 |
@@
Visual Enhancing Decoding Strategy.
Although LVLMs can accurately recognize visual elements, LVLMs have difficulty fully interpreting those elements in the context of the input cue and effectively linking that recognition to their internal knowledge. We follow Visual Description Grounded Decoding (VDGD) (Ghosh et al., 2024) by first generating a detailed description of the image and appending it as a prefix to the instruction. The prompt template is adopted from (Ghosh et al., 2024):
<image> I have been given this image to complete the task described as: inst. To help me complete the task, describe the given image in detail. In the case of real-world scenes, please include all foreground and background objects in the description, their properties (like color, shape, etc.), their relations with other objects, their count, and all other components in the image. In case of non-real-world scenes, like charts, graphs, tables, etc., please describe the table, mention all numbers (if any), mention the written text, and all other details. Experiments are performed on hallucination evaluation benchmarks( i.e., POPE and CHAIR) and the general ability benchmark (i.e., MMbench). We re-implement VDGD based on official codes on LLaVA-1.5 7B.
Table 15 demonstrates the effectiveness of VDGD (Ghosh et al., 2024) in LVLM’s hallucination alleviation and general reasoning ability. However, the grounding visual descriptions, generated by LVLMs themself, enhance the visual perception reasoning capabilities while might inevitably contain hallucinations. Therefore, VDGD is inferior in the POPE (adversial) subset, which prioritizes co-occurring objects which are not present in the image. Meanwhile,
VDGD shares somewhat similar motivations in enhancing vision information via Eq. 6 as we analyzed.
The experiments in Table 2 are consistent with the above results, indicating that boosting the vision information is effective in mitigating hallucinations but is less effective in complex (i.e., adversarial) environments.
CHAIRs | CHAIRi | MMbench | |||||
| Greedy | 88.8 | 79.1 | 49.6 | 14.4 | 64.4 | ||||
| OPERA | 88.9 | 79.2 | 45.2 | 12.7 | 64.4 | ||||
| VDGD⋆ | 89.0 | 79.4 | 46.7 | 13.7 | 65.2 | ||||
| SID⋆ | 89.3 | 83.3 | 44.2 | 12.2 | 65.0 |